# Thermal description of braneworld effective theories

### v1

@@ Abstract
What exactly does radion stabilization do to the critical curves in the $(\Theta,{\cal KT})$ plot? We discussed that radion settles to a stable minimum of the radion potential with any residue dynamics being fluctuations about this stable minimum. As previously mentioned, the on-brane theory would eventually converge to GR. Thus, modulus stabilization effectively does two jobs - first, it straightens the critical curves - as the slopes depend on Brans-Dicke functions which in turn only depend on the radion field; they become almost constant - and second, it takes all possible cosmolgical solutions to the lower half of the corresponding critical line, ensuring the convergence of the theory to GR (see Fig.1). One may wonder that upon including a stabilizing field, we would get the radion potential contribution on the right side of Eq.(18) which would hinder the critical curves from being straight lines. But since the dynamics is now restricted to around the stable minimum, we can take the radion potential to be effectively $V(\Phi_{\pm})\propto\Phi_{\pm}^{2}$. This makes the contribution $(V^{\prime}-\frac{2V}{\Phi_{\pm}})$ vanish and hence, we still retain the form of Eq.(18).
@@
This essentially tells us that the $\cal KT$} dynamics on either brane is equivalent upto a conformal factor - if we know the dynamics on one brane, we know the dynamics on the other. At a glance, this may seem like a non-local effect, but it is truly an artefact of the facet that the Brans-Dicke field on either brane depends on the radion field, which governs the inter-brane separation. Even in the general scenario, one could do the conformal transformation explicitly to get the dynamics on TeV brane but then the relation would not be this direct.
@@
Thus, if the expansion scalar for a given geometry obeys the above inequality for a given $f(R)$ theory, then the gravity theory necessarily diverges away from GR. We can analyze the $V^{\prime}-\frac{2V}{\Psi}$ combination further in the Jordan frame but to obtain a consistent criterion on the underlying f(R) theory (to distinguish whether it heats or cools gravity), we have to impose non-trivial constraints on the theory which are not very well physically motivated. So, for the case at hand, we instead switch over to the Einstein frame through a conformal transformation [60] which makes us rely on only the constraints which ensure stability of the f(R) theory i.e . $f^{\prime}(R)>0,f^{\prime\prime}(R)>0$ [61, 62]. But, then we shall focus on the chemical potential dynamics instead of temperature (which is trivially zero in this frame) [31],
@@
where $\square\Phi=V^{\prime}(\Phi)$ and $\Phi$ is the Einstein frame (minimally coupled) scalar. $V^{\prime}(\Phi)$ turns out to be $\frac{1}{\sqrt{3}}[Rf^{\prime}(R)-2f(R)]$. Thus, from the prior equation, we see that $Rf^{\prime}(R)-2f(R)<0$ to help the theory tend towards GR, which essentially boils down to $\frac{d}{dR}\ln(\frac{f(R)}{R^{2}})<0$. In other words, $f(R)$ has to grow slower than $R^{2}$ to aid the attractor mechanism. For a Starobinsky theory $f(R)=R+\gamma R^{2}$, this attractor-to-GR condition reduces to having positive curvature theories ($R>0$). We should also notice that pure quadratic gravity $f(R)=R^{2}$ neither helps nor hinders the gravity theory to converge to GR.
@@
Discussion— We have provided a consistent thermal description of scalar-tensor effective theories in braneworld scenarios emphasizing the attractor mechanism towards GR. We explored the two-brane Randall-Sundrum model where the low-energy theory on either brane is of scalar-tensor nature with the extra-dimensional radion playing the role of the Brans-Dicke scalar, and found that the negative tension TeV brane acts as a better “cooler” than the positive tension Planck brane in the sense that the TeV brane allows a much larger class of solutions to tend towards GR - this could also be seen when we confined to an on-brane FLRW metric. We further explored the detunings of the brane tensions in context of a non-vanishing gravitational contribution to radion potential. We found that in the “almost flat Planck brane” and the “almost flat TeV brane” regimes, identical detuning of the brane tensions ensures identical signature for the radion potential contribution to the $\cal KT$} dynamics of the gravity theory (whether it “cools” or “heats” gravity) on either brane. We also identified a case where simultaneous detuning of the brane tensions could result in a “cooling” contribution without adhering to a regime if the Planck brane geometry is of the anti-de Sitter type. We then discussed that the radion potential contribution by a bulk Goldberger-Wise scalar in fact “cools” down gravity for a positive mass squared scalar (which is known to stabilize the radion) and “heats” gravity for a negative mass squared scalar (which fails to stabilize the radion). We highlighted that radion stabilization smoothens the critical curves to straight lines and considering residue fluctuations about the stabilized radion, we showed that the $\cal KT$} dynamics on either brane is equivalent upto a conformal factor.
@@
One should note that our present analysis is by no means exhaustive, and that there may be several interesting extensions - RS with a time-dependent stabilizing field [63], the non-flat brane RS model [10, 58], $f(R)$ braneworlds [59, 64], multi-brane extensions [65, 66] and other higher-dimensional modified gravity scenarios where the low-energy effective theory is either scalar-tensor or more generally, satisfies the Eckart-Fourier constitutive equation [46]. In this context, it would also be interesting to explore braneworld models like the Dvali-Gabadadze-Porrati (DGP) model [67] and cascading gravity [68] whose effective theories are Horndeski gravity, which is known to have a first-order thermodynamic interpretation. Some of these will also be addressed in our future works.
@@
Acknowledgement— SB is supported through INSPIRE-SHE Scholarship by the Department of Science and Technology, Government of India.

### v2

@@ Abstract
What exactly does radion stabilization do to the critical curves in the $(\Theta,{\cal KT})$ plot? We discussed that radion settles to a stable minimum of the radion potential with any residue dynamics being fluctuations about this stable minimum. As previously mentioned, the on-brane theory would eventually converge to GR. Thus, modulus stabilization effectively does two jobs - first, it straightens the critical curves - as the slopes depend on Brans-Dicke functions which in turn only depend on the radion field; they become almost constant - and second, it takes all possible cosmolgical solutions to the lower half of the corresponding critical line, ensuring the convergence of the theory to GR (see Fig.1). One may wonder that upon including a stabilizing field, we would get the radion potential contribution on the right side of Eq.(16) which would hinder the critical curves from being straight lines. But since the dynamics is now restricted to around the stable minimum, the radion potential is effectively $V(\Phi_{\pm})\propto\Phi_{\pm}^{2}$. This makes the contribution $(V^{\prime}-\frac{2V}{\Phi_{\pm}})$ vanish and hence, we still retain the form of Eq.(16).
@@
This essentially tells us that the $\cal KT$} dynamics on either brane is equivalent upto a conformal factor - if we know the dynamics on one brane, we know the dynamics on the other. At a glance, this may seem like a non-local effect, but it is truly an artefact of the facet that the Brans-Dicke field on either brane depends on the radion field, which governs the inter-brane separation. In the general scenario, one could do the conformal transformation explicitly to get the dynamics on TeV brane but then the relation would not be this direct.
@@
Thus, if the expansion scalar for a given geometry obeys the above inequality for a given $f(R)$ theory, then the gravity theory necessarily diverges away from GR. We can analyze the $V^{\prime}-\frac{2V}{\Psi}$ combination further in the Jordan frame but to obtain a consistent criterion on the underlying f(R) theory (to distinguish whether it heats or cools gravity), we have to impose non-trivial constraints on the theory which are not very well physically motivated. So, for the case at hand, we instead switch over to the Einstein frame through a conformal transformation [61] which makes us rely on only the constraints which ensure stability of the f(R) theory i.e . $f^{\prime}(R)>0,f^{\prime\prime}(R)>0$ [37, 13]. But, then we shall focus on the chemical potential dynamics instead of temperature (which is trivially zero in this frame) [63, 30],
@@
where $\mu=\sqrt{-\nabla^{a}\Phi\nabla_{a}\Phi}$, $\square\Phi=V^{\prime}(\Phi)$ and $\Phi$ is the Einstein frame (minimally coupled) scalar. $V^{\prime}(\Phi)$ turns out to be $\frac{1}{\sqrt{3}}[Rf^{\prime}(R)-2f(R)]$. Thus, from the prior equation, we see that $Rf^{\prime}(R)-2f(R)<0$ to help the theory tend towards GR, which essentially boils down to $\frac{d}{dR}\ln(\frac{f(R)}{R^{2}})<0$. In other words, $f(R)$ has to grow slower than $R^{2}$ to aid the attractor mechanism. For a Starobinsky theory $f(R)=R+\gamma R^{2}$, this attractor-to-GR condition reduces to having positive curvature theories ($R>0$). We should also notice that pure quadratic gravity $f(R)\propto R^{2}$ neither helps nor hinders the gravity theory to converge to GR.
@@
Discussion— We have provided a consistent thermal description of scalar-tensor effective theories in braneworld scenarios emphasizing the attractor mechanism towards GR. We explored the two-brane Randall-Sundrum model where the low-energy theory on either brane is of scalar-tensor nature with the extra-dimensional radion playing the role of the Brans-Dicke scalar. We further explored the detunings of the brane tensions in context of a non-vanishing gravitational contribution to radion potential, and found that in the “almost flat Planck brane” and the “almost flat TeV brane” regimes, non-dynamical detunings of the brane tensions exactly determine the signature for the radion potential contribution to the $\cal KT$} dynamics of the gravity theory (whether it “cools” or “heats” gravity) on either brane. We also identified a case where simultaneous detuning of the brane tensions could result in a “cooling”(“heating”) contribution without adhering to a regime if the Planck brane geometry is of the anti-de Sitter (de Sitter) type. We then discussed that the radion potential contribution by a bulk Goldberger-Wise scalar in fact “cools” down gravity for a positive mass squared scalar (which is known to stabilize the radion) and “heats” gravity for a negative mass squared scalar (which fails to stabilize the radion). Restricting to fluctuations about the stabilized radion (upon including a bulk stabilizing field), we showed that the critical curves get smoothened to straight lines and that the $\cal KT$} dynamics on either brane is equivalent upto a conformal factor.
@@
One should note that our present analysis is by no means exhaustive, and that there may be several interesting extensions - RS with a time-dependent stabilizing field [14], the non-flat brane RS model [24, 4], $f(R)$ braneworlds [6, 27], multi-brane extensions [16, 11] and other higher-dimensional modified gravity scenarios where the low-energy effective theory is either scalar-tensor or more generally, satisfies the Eckart-Fourier constitutive equation [26].
In this context, it would also be interesting to explore braneworld models like the Dvali-Gabadadze-Porrati (DGP) model [60] whose effective theories are of the general Horndeski variety, which is known to have a first-order thermodynamic interpretation. Some of these will also be addressed in our future works.
@@
Acknowledgement— SB is supported through INSPIRE-SHE Scholarship by the Department of Science and Technology, Government of India. SB acknowledges helpful discussions with participants of the J.V. Narlikar Memorial Conference on Cosmology and Astrophysics, 2025 held at Visva-Bharati, Santiniketan.